OpenStudy (anonymous):

How do I find the current if the problem only gives me resistance and power?

OpenStudy (joannablackwelder):

P=IV and V=IR so P=I(IR)

OpenStudy (joannablackwelder):

Thus I=sqrt(P/R)

OpenStudy (anonymous):

So an example being, A device has a resistance of 67 ohms and a Power rating of 43 watts what is the current? How would you do this?

OpenStudy (anonymous):

I am still confused about how you got to that formula.

OpenStudy (joannablackwelder):

Are you familiar with the two formluas I mentioned before substituting?

OpenStudy (anonymous):

Yes

OpenStudy (joannablackwelder):

43 W=I^2 * 67 ohms, solve for I.

OpenStudy (joannablackwelder):

Using dimensional analysis, I get the unit of A.

OpenStudy (anonymous):

You are going wayyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyy beyond the 9th grade level with vocab.

OpenStudy (joannablackwelder):

Do you mean dimensional analysis? That is just ignoring the numbers and looking only at the units.

OpenStudy (anonymous):

OH. I am so stupid. o. o

OpenStudy (anonymous):

What is the unit "A"

OpenStudy (joannablackwelder):

No worries. A is the unit of current.

OpenStudy (anonymous):

Power consumed by heating = P = I^2 R = current squared times resistance so I = sqrt(P/R), as noted by J-B-.